12. Consider the region bounded above by the function ?=1/(?+2)2(?+6)^2 and below by the xy-plane for x≥0 and ?≥0.

(1 point) Consider the region bounded above by the function z = - 2 (x + 2)2(y + 62 an and below by the xy-plane for x > 0
Answer #1

+ 1² (4 +62 x 30, y = 0 & xy plane Shadow is entire the 1st Quadrant od volume = JJ z dy da a cled 3 volunne = f (x +27 (y +65 23°Cyr 6) = dx dy Jo (in dy 29+0= dy (2tyrb)], - Volume is Ž me